# Large-scale Pre-training for Grounded Video Caption Generation

### v2

@@ Abstract
We propose a novel approach for captioning and object grounding in video, where the objects in the caption are grounded in the video via temporally dense bounding boxes. We introduce the following contributions. First, we present a large-scale automatic annotation method that aggregates captions grounded with bounding boxes across individual frames into temporally dense and consistent bounding box annotations. We apply this approach on the HowTo100M dataset to construct a large-scale pre-training dataset, named HowToGround1M. We also introduce a Grounded Video Caption Generation model, dubbed GROVE, and pre-train the model on HowToGround1M. Second, we introduce a new dataset, called iGround, of 3500 videos with manually annotated captions and dense spatio-temporally grounded bounding boxes. This allows us to measure progress on this challenging problem, as well as to fine-tune our model on this small-scale but high-quality data. Third, we demonstrate that our approach achieves state-of-the-art results on the proposed iGround dataset compared to a number of baselines, as well as on the VidSTG and ActivityNet-Entities datasets. We perform extensive ablations that demonstrate the importance of pre-training using our automatically annotated HowToGround1M dataset followed by fine-tuning on the manually annotated iGround dataset and validate the key technical contributions of our model.
@@ 1 Introduction
In this work, we address this key limitation by the following three contributions. First, to address the issue of limited training data, we introduce a large-scale automatic annotation method leveraging an existing model for grounded still-image captioning together with an LLM to summarize frame-level captions into video-level captions. The LLM is also tasked to perform temporally consistent bounding box annotation, associating frame-level phrases that correspond to objects with video-level phrases, resulting in a video-level caption grounded with multiple object tubes with consistent natural language labels. We apply this approach to videos from the HowTo100M [22] dataset, which results in a new large-scale pre-training dataset, namely HowToGround1M, of more than 1M videos for this problem. This automatic annotation method is coupled with a proposed GROunded Video Caption gEneration model, called GROVE. The key technical contributions of this model include: (i) spatio-temporal adapters, which enable efficient modeling of spatio-temporal information in video; (ii) a bounding box decoder that outputs temporally coherent bounding boxes in video and (iii) a temporal objectness head that explicitly models objects that temporary leave the frame or are occluded. We pre-train the GROVE model on the proposed large-scale automatically annotated HowToGround1M dataset. Second, we introduce a new manually annotated dataset for the grounded caption generation task, which we name iGround. The dataset contains 3500 videos and more than 230,000 annotated object bounding boxes. We split this dataset into train/val/test sets (2000/500/1000, respectively). This allows us to measure progress on this challenging problem, as well as to fine-tune our model on small-scale but high-quality data. Third, our results demonstrate that our GROVE model achieves state-of-the-art performance on the proposed iGround dataset as well as on the VidSTG [43] and ActivityNet-Entities [45] datasets. We perform extensive ablations that demonstrate the importance of pre-training using our automatically annotated HowToGround1M dataset followed by fine-tuning on the manually annotated iGround dataset and validate the key technical contributions of our model.
@@ 2 Related Work
Dataset Annot.
Multiple
Multi-object
Num.
Num.
@@
Datasets for spatio-temporal grounding in video. The existing datasets [34, 43, 11, 33, 4, 45] for spatio-temporal video grounding are relatively small scale as they rely on manual annotation, which is tedious and time consuming. Typically, the existing datasets also focus on localizing a single spatio-temporal tube for each short textual description  [34, 43, 33, 4], which can be a limiting factor in instructional videos where multiple objects are often manipulated. In contrast, both our automatically annotated HowToGround1M dataset as well as our manually-labelled iGround dataset contain multiple spatio-temporal bounding boxes grounding multiple objects described in the caption. We compare our HowToGround1M dataset as well as our iGround dataset with exisiting video grounding datasets in Table 2. ActivityNet-Entities [45] is closest to ours in that it provides grounded captions – multiple objects per frame for annotated noun phrases in the caption. Nevertheless, the authors annotate a single frame per object in the video segment while both our automatically annotated HowToGround1M as well as our manually annotated iGround dataset have densely annotated frames per video segment. Moreover, HowToGround1M has the largest number of videos across all datasets and the largest number of annotated instances. More statistics and analysis of our proposed datasets is provided in the supplementary material.
Spatio-temporal grounding in video. Spatio-temporal grounding [34, 43, 39, 33, 17, 4, 32, 42, 13, 10, 38] aims to predict a single spatio-temporal tube enclosing an event described in a natural language query. Early approaches relied on object detection features [43, 34]. [43] introduced a spatio-temporal graph encoder, while [34] and [39] adapted transformers for multi-modal grounding. Some works employed contrastive learning on large-scale instructional videos, leveraging weak supervision from HowTo100M [33, 4]. Others explored architectural innovations, such as two-stream models for appearance and motion [17] or context-guided decoding for object-centric grounding [10]. However, these methods lack a text generation component, as they assume the query is given, and they cannot generate multiple spatio-temporal tubes for multiple objects. Our approach addresses this issue with the GROVE model, a novel automatic annotation method, and a manually-labeled dataset for fine-tuning and evaluation. Closest to our work is [45], which separately performs captioning and grounding. In contrast, our GROVE model jointly (i) captions, (ii) tags groundable noun phrases, and (iii) grounds them with temporally dense bounding boxes while handling occlusions. Experimental results demonstrate that we outperform [45] by a margin.
@@ 3.1 Automatic annotation method
Stage 2: Video-level caption aggregation. The previous stage provides outputs for individual frames, which are, however, not temporally consistent. We address this issue in Stage 2. From the frame-level captions, we generate a video-level caption that highlights the most salient actions and objects while ensuring consistency in noun phrase annotations across frames. We achieve this by prompting Llama-2 [35], as described next. Since the grounded captioning model produces lengthy captions with extraneous details, we first extract Subject-Verb-Object (SVO) triplets and relevant adpositional phrases using Part-of-Speech (POS) tagging. These structured triplets serve as input to the LLM, encapsulating subjects, actions, and relevant objects. To enhance accuracy, we perform in-context learning by providing example pairs of frame-level SVO triplets and their corresponding video-level captions (see supplementary material for details). Given a new set of SVO triplets, the LLM generates a video-level caption and tags noun phrases corresponding to objects of interest within <p></p> tags. Figure 2 (middle) illustrates this process. In experiments, we compare the above approach with providing the LLM with full captions from Stage 1 instead of extracting Subject-Verb-Object triplets from the caption to assess the impact of additional context.
Stage 3: Temporally consistent bounding box annotation. While Stage 2 ensures a consistent video-level caption, noun phrase annotations remain inconsistent across frames due to the use of an image-based model. To resolve this, we introduce temporal labeling of objects, ensuring that bounding boxes corresponding to the same object are consistently labeled throughout the video using the video-level noun phrases. These consistently labeled bounding boxes form video object tracks. We formulate this as a text classification task and prompt the LLM with in-context learning. The input consists of frame-level noun phrases to be classified and the video-level noun phrases serving as class labels. The supplementary material provides details on the prompt. Figure 2 (right) illustrates this process. An alternative approach is to use a visual tracker, such as [14], for associating bounding boxes over time, which we compare against in Sec. 6.2.
Upon completing all three stages, we obtain videos with automatically generated captions, grounded bounding boxes, and temporally coherent labels aligned with the noun phrases in the caption. Additional details are provided in the supplementary material.
@@ 3.2 The HowToGround1M dataset
We apply our automatic annotation method to Internet instructional videos from the HowTo100M dataset [22]. We choose HowTo100M due to its diversity of actions, scenes, objects and lighting conditions. We apply our approach to clips obtained using time stamps from the HowToCaption [31] version of the dataset. These clips contain meaningful events in the video identified based on LLM analysis of the associated narrations. In detail, we randomly sample 1M video clips from HowTo100M videos using start/end timestamps from HowToCaption. The videos from HowTo100M have variable frame rates usually ranging in 25-30 fps, and we process the videos at 5fps. The majority of the clips are 8 seconds long with a spatial resolution of 455$\times$256 pixels. We run our automatic annotation method on this set of data to obtain our HowToGround1M pre-training dataset. The resulting dataset contains 1M videos, with 1M captions containing 3.2M noun phrases. The captions contain 18.6k unique terms and 142k unique noun phrases. Overall, the dataset contains 43.6M annotated frames with 80.1M bounding boxes. This dataset is used to pre-train the GROVE model, described next.
@@ 4 The GROVE Model
Figure 3: An overview of our GROVE model for grounded video caption generation. Dashed red rectangles outline the key technical contributions enabling grounded caption generation in video and include: (i) spatio-temporal adapters; (ii) the bounding box decoder and (iii) the temporal objectness head.
We introduce a GROunded Video caption gEneration model, called GROVE, see Figure 3. The input to the model is a video clip with $T$ frames (left) and the output is a natural language caption (right) together with $N$ spatio-temporal bounding boxes localizing the individual noun phrases in the video together with a objectness score indicating the presence / absence of the object in a specific frame (top). We build on the GLaMM [28] model, which is a state-of-the-art method for still image-based grounded caption generation, and extend it to the video domain. The key technical components enabling grounded caption generation in videos are (shown in dashed red rectangles in Figure 3): i) the spatio-temporal adapters with pooling which enable modelling temporal information efficiently; ii) the bounding box decoder which allows re-using large-scale pretrained decoder weights [28]; and iii) the temporal objectness head for modelling objects that temporary leave the frame or are occluded. Details are given below. We assess the importance of these components in Section 6. Additional details of the GROVE model and the training procedure are in the supplementary material.
Spatio-temporal adapters and pooling. The visual information is encoded using the Global Video Encoder $\mathcal{V}_{e}(\cdot)$, which represents the video globally for captioning, and the Grounding Video Encoder $\mathcal{V}_{g}(\cdot)$, which represents the fine-grained details for grounding. We build these encoders by adapting the respective pre-trained image encoders [28]. We achieve this by interleaving spatio-temporal adapter layers (denoted as $a()$) between the image-based encoder layers. To stabilise training, we add residual connections and introduce a learnable parameter that is multiplied by the adapter’s output and starts from 0 at the beginning of training [1]. By doing so, at the beginning of training the adapter’s output is effectively cancelled out and the network observes only the original encoder’s output. As training progresses, the learnable parameter is tuned and the network automatically adjusts the contribution of the adapter based on the gradients of the loss. In detail, the adapter layer performs $a(o)=o+tanh(\alpha)\times f(o)$, where $o$ is the output of the preceding encoder layer, $\alpha$ is the tunable parameter that is initialised to 0 and passes through a $tanh$ activation and $f(\cdot)$ is the adapter layer. As feeding the full video tokens $o_{e}$ to the LLM is computationally prohibitive, we introduce a spatio-temporal pooling function after the output of $\mathcal{V}_{e}(\cdot)$, i.e., $o_{p}=p(o_{e})$.
@@ 5 Manually annotated iGround dataset
For the iGround dataset we select 3500 clips from the HowTo100M dataset [22] by sampling ‘interesting’ videos that typically include dynamic events or actions that are clear and distinguishable. In those events/actions, people usually interact with objects. We make sure that the 3500 clips do not overlap with those in the HowToGround1M dataset described in section 3. We split the set into 2000/500/1000 video clips to form the train/val/test sets, respectively. The video annotation itself consists of 3 steps. The first step entails watching the video and providing a natural language description of what is happening in the video and the objects that are being manipulated. Note, that we are interested in the active objects, i.e. objects that humans interact with, rather than densely describing all objects in the scene. In the second step, bounding boxes are annotated for all visible instances of humans/objects mentioned in the caption that has been provided in the previous step. Finally, each bounding box is annotated with a short phrase or word that should match exactly a short phrase/word from the caption. More information about the annotation procedure including the exact annotation guidelines for the raters and the mechanisms for assuring the consistency and the overall quality of the rater annotations, can be found in the supplementary material.
@@ 6.1 Datasets and evaluation metrics
ActivityNet-Entities. We follow [45] and report F1all, F1all_per_sent, F1loc, and F1loc_per_sent. In F1all, a region prediction is considered correct if the associated noun phrase is both correctly predicted (exact match) and correctly localised (IoU $>$ 0.5). F1loc considers only localisation accuracy ignoring errors in the generated noun phrases. In these metrics, accuracies are averaged across noun categories while in F1all_per_sent and F1loc_per_sent accuracies are averaged across sentences.
@@
Implementation details. All implementation details including architectural choices of the GROVE model as well as training and inference details can be found in the supp. mat.
@@
b. GROVE - PT (Ours) 14.5 49.9 26.7 23.0
c. GROVE - PT+FT (Ours) 21.7 85.4 31.9 25.4
@@
e. GROVE - PT (Ours) 14.5 49.9 34.2 25.4
f. GROVE - FT (Ours) 21.0 81.7 15.7 17.4
g. GROVE - PT+FT (Ours) 21.7 85.4 40.8 28.6
@@
GROVE PT+FT (Ours) 62.9
Table 3: State-of-the-art comparison of spatial grounding on the VidSTG [43] test set. All models use ground truth temporal localization. Large-scale pretraining (PT+FT) results in an improvement over fine-tuning only (FT) for our model GROVE.

### v3

@@ Abstract
We propose a novel approach for captioning and object grounding in video, where the objects in the caption are grounded in the video via temporally dense bounding boxes. We introduce the following contributions. First, we present a large-scale automatic annotation method that aggregates frame-level captions grounded with bounding boxes into temporally dense and consistent annotations. We apply this approach on the HowTo100M dataset to construct a large-scale pre-training dataset, named HowToGround1M. We also introduce a Grounded Video Caption Generation model, dubbed GROVE, and pre-train the model on HowToGround1M. Second, we introduce iGround–a dataset of 3513 videos with manually annotated captions and dense spatio-temporally grounded bounding boxes. This allows us to measure progress on this challenging problem, as well as to fine-tune our model on this small-scale but high-quality data. Third, we demonstrate that our approach achieves state-of-the-art results on the proposed iGround dataset, as well as on the VidSTG, ActivityNet-Entities, GroundingYouTube, and YouCook-Interactions datasets. Our ablations demonstrate the importance of pre-training on our automatically annotated HowToGround1M dataset followed by fine-tuning on the manually annotated iGround dataset and validate the key technical contributions of our model. The dataset and code are available at https://ekazakos.github.io/grounded_video_caption_generation/.
@@ 1 Introduction
In this work, we address this key limitation by the following three contributions. First, to address the issue of limited training data, we introduce a large-scale automatic annotation method leveraging an existing model for grounded still-image captioning together with an LLM to summarize frame-level captions into video-level captions. The LLM is also tasked to perform temporally consistent bounding box annotation, associating frame-level phrases that correspond to objects with video-level phrases, resulting in a video-level caption grounded with multiple object tubes with consistent natural language labels. We apply this approach to videos from the HowTo100M [23] dataset, which results in a new large-scale pre-training dataset, namely HowToGround1M, of 1M videos for this problem. This automatic annotation method is coupled with a proposed GROunded Video Caption gEneration model, called GROVE. The key technical contributions of this model include: (i) spatio-temporal adapters, which enable efficient modeling of spatio-temporal information in video; (ii) a bounding box decoder that outputs temporally coherent bounding boxes in video and (iii) a temporal objectness head that explicitly models objects that temporary leave the frame or are occluded. We pre-train the GROVE model on the proposed large-scale automatically annotated HowToGround1M dataset. Second, we introduce a new manually annotated dataset for the grounded caption generation task, which we name iGround. The dataset contains 3513 videos and more than 230,000 annotated object bounding boxes. We split this dataset into train/val/test sets (2013/500/1000, respectively). This allows us to measure progress on this challenging problem, as well as to fine-tune our model on small-scale but high-quality data. Third, our results demonstrate that our GROVE model achieves state-of-the-art performance on the proposed iGround dataset, as well as on the VidSTG [48], ActivityNet-Entities [50], GroundingYouTube [4], and YouCook-Interactions [37] datasets. We perform extensive ablations that demonstrate the importance of pre-training using our automatically annotated HowToGround1M dataset followed by fine-tuning on the manually annotated iGround dataset and validate the key technical contributions of our model. For additional details, see the appendix.
@@ 2 Related Work
Dataset
@@
Datasets for spatio-temporal grounding in video. The existing datasets [38, 48, 12, 37, 5, 50] for spatio-temporal video grounding are relatively small scale as they rely on manual annotation, which is tedious and time consuming. Typically, the existing datasets also focus on localizing a single spatio-temporal tube for each short textual description  [38, 48, 37, 5], which can be a limiting factor in instructional videos where multiple objects are often manipulated. In contrast, both our automatically annotated HowToGround1M dataset as well as our manually-labelled iGround dataset contain multiple spatio-temporal bounding boxes grounding multiple objects described in the caption. We compare our HowToGround1M dataset as well as our iGround dataset with exisiting video grounding datasets in Table 2. ActivityNet-Entities [50] is closest to ours in that it provides grounded captions – multiple objects per frame for annotated noun phrases in the caption. Nevertheless, the authors annotate a single frame per object in the video segment while both our automatically annotated HowToGround1M as well as our manually annotated iGround dataset have densely annotated frames per video segment. Moreover, HowToGround1M has the largest number of videos across all datasets and the largest number of annotated instances. More statistics and analysis of our proposed datasets is provided in appendix D.
Spatio-temporal grounding in video. Spatio-temporal grounding [38, 48, 43, 37, 18, 5, 36, 47, 14, 11, 42] aims to predict a single spatio-temporal tube enclosing an event described in a natural language query. Early approaches relied on object detection features [48, 38]. [48] introduced a spatio-temporal graph encoder, while [38] and [43] adapted transformers for multi-modal grounding. Some works employed contrastive learning on large-scale instructional videos, leveraging weak supervision from HowTo100M [37, 5]. Others explored architectural innovations, such as two-stream models for appearance and motion [18] or context-guided decoding for object-centric grounding [11]. However, these methods lack a text generation component, as they assume the query is given, and they cannot generate multiple spatio-temporal tubes for multiple objects. Our approach addresses this issue with the GROVE model, a novel automatic annotation method, and a manually-labeled dataset for fine-tuning and evaluation. More pertinent to our work is [50], which separately performs captioning and grounding. In contrast, our GROVE model jointly (i) captions and (ii) grounds noun phrases with temporally dense bounding boxes while handling occlusions. Concurrent to our work is [26], which targets mask-level video grounding and trains on semi-automatic annotations from existing benchmarks. In contrast, we focus on bounding boxes, pre-train with large-scale fully automatic annotations and fine-tune on human-annotated data. We outperform both [26] and [50] .
@@ 3 Large-scale generation of grounded captions
Figure 3: An overview of our GROVE model for grounded video caption generation. Dashed red rectangles outline the key technical contributions enabling grounded caption generation in video and include: (i) spatio-temporal adapters; (ii) the bounding box decoder and (iii) the temporal objectness head. VL: vision-to-language projection, LQ: language-to-query projection.
@@ 3.1 Automatic annotation method
Stage 2: Video-level caption aggregation. The previous stage provides outputs for individual frames, which are, however, not temporally consistent. We address this issue in Stage 2. From the frame-level captions, we generate a video-level caption that highlights the most salient actions and objects while ensuring consistency in noun phrase annotations across frames. We achieve this by prompting Llama-2 [39], as described next. Since the grounded captioning model produces lengthy captions with extraneous details, we first extract Subject-Verb-Object (SVO) triplets and relevant adpositional phrases using Part-of-Speech (POS) tagging. These structured triplets serve as input to the LLM, encapsulating subjects, actions, and relevant objects. To enhance accuracy, we perform in-context learning by providing example pairs of frame-level SVO triplets and their corresponding video-level captions (see Figure 11 for details). Given a new set of SVO triplets, the LLM generates a video-level caption and tags noun phrases corresponding to objects of interest within <p></p> tags. Figure 2 (middle) illustrates this process. In experiments, we compare the above approach with providing the LLM with full captions from Stage 1 instead of extracting Subject-Verb-Object triplets from the caption to assess the impact of additional context.
Stage 3: Temporally consistent bounding box annotation. While Stage 2 ensures a consistent video-level caption, noun phrase annotations remain inconsistent across frames due to the use of an image-based model. To resolve this, we introduce temporal labeling of objects, ensuring that bounding boxes corresponding to the same object are consistently labeled throughout the video using the video-level noun phrases. These consistently labeled bounding boxes form video object tracks. We formulate this as a text classification task and prompt the LLM with in-context learning. The input consists of frame-level noun phrases to be classified and the video-level noun phrases serving as class labels. Figure 12 provides details on the prompt. Figure 2 (right) illustrates this process. An alternative approach is to use a visual tracker, such as [15], for associating bounding boxes over time, which we compare against in Sec. 6.2.
Upon completing all three stages, we obtain videos with automatically generated captions, grounded bounding boxes, and temporally coherent labels aligned with the noun phrases in the caption. Additional details are provided in appendix F.
@@ 3.2 The HowToGround1M dataset
[…]
@@ 4 The GROVE Model
We introduce a GROunded Video caption gEneration model, called GROVE, see Figure 3. The input to the model is a video clip with $T$ frames (left) and the output is a natural language caption (right) together with $N$ spatio-temporal bounding boxes localizing the individual noun phrases in the video together with a objectness score indicating the presence / absence of the object in a specific frame (top). We build on the GLaMM [31] model, which is a state-of-the-art method for still image-based grounded caption generation, and extend it to the video domain. The key technical components enabling grounded caption generation in videos are (shown in dashed red rectangles in Figure 3): i) the spatio-temporal adapters with pooling which enable modelling temporal information efficiently; ii) the bounding box decoder which allows re-using large-scale pretrained decoder weights [31]; and iii) the temporal objectness head for modelling objects that temporary leave the frame or are occluded. Details are given below. We assess the importance of these components in Section 6. Additional details of the GROVE model and the training procedure are in appendix E.
Spatio-temporal adapters and pooling. The visual information is encoded using the Global Video Encoder $\mathcal{V}_{e}(\cdot)$, which represents the video globally for captioning, and the Grounding Video Encoder $\mathcal{V}_{g}(\cdot)$, which represents the fine-grained details for grounding. We build these encoders by adapting the respective pre-trained image encoders [31]. We achieve this by interleaving spatio-temporal adapter layers (denoted as $a()$) between the image-based encoder layers. To stabilise training, we add residual connections and introduce a learnable parameter that is multiplied by the adapter’s output and starts from 0 at the beginning of training [1]. By doing so, at the beginning of training the adapter’s output is effectively cancelled out and the network observes only the original encoder’s output. As training progresses, the learnable parameter is tuned and the network automatically adjusts the contribution of the adapter based on the gradients of the loss. In detail, the adapter layer performs $a(o)=o+tanh(\alpha)\times f(o)$, where $o$ is the output of the preceding encoder layer, $\alpha$ is the tunable parameter that is initialised to 0 and passes through a tanhtanh activation and $f(\cdot)$ is the adapter layer. As feeding the full video tokens $o_{e}$ to the LLM is computationally prohibitive, we introduce a spatio-temporal pooling function after the output of $\mathcal{V}_{e}(\cdot)$, i.e., $o_{p}=p(o_{e})$.
@@ 5 Manually annotated iGround dataset
For the iGround dataset, we select 3513 clips from the HowTo100M dataset [23] by sampling ‘interesting’ videos that typically include dynamic events or actions that are clear and distinguishable. In those events/actions, people usually interact with objects. We make sure that the 3513 clips do not overlap with those in the HowToGround1M dataset described in section 3. Moreover, we ensure that there is no video ID overlap with any of the downstream datasets. We split the set into a training set (2013 clips), a validation set (500 clips), and a test set (1000 clips), ensuring no overlap in video IDs between them. The video annotation itself consists of 3 steps. The first step entails watching the video and providing a natural language description of what is happening in the video and the objects that are being manipulated. Note, that we are interested in the active objects, i.e. objects that humans interact with, rather than densely describing all objects in the scene. In the second step, bounding boxes are annotated for all visible instances of humans/objects mentioned in the caption that has been provided in the previous step. Finally, each bounding box is annotated with a short phrase or word that should match exactly a short phrase/word from the caption. More information on the annotation procedure, including annotation guidelines for the raters and mechanisms to ensure the consistency and overall quality of the raters’ annotations, can be found in appendix G.
@@ 6.1 Datasets and evaluation metrics
ActivityNet-Entities. We follow [50] and report F1all, F1all_per_sent, F1loc, and F1loc_per_sent. In F1all, a region prediction is considered correct if the associated noun phrase is both correctly predicted (exact match) and correctly localised (IoU >> 0.5). F1loc considers only localisation accuracy ignoring errors in the generated noun phrases. In these metrics, accuracies are averaged across noun categories while in F1all_per_sent and F1loc_per_sent accuracies are averaged across sentences.
@@
Implementation details. All implementation details including architectural choices of the GROVE model as well as training and inference details can be found in appendix E.
@@
b. GROVE - PT (Ours) 14.3 50.6 27.0 22.5
c. GROVE - PT+FT (Ours) 21.4 83.5 31.7 26.2
@@
e. GROVE - PT (Ours) 14.3 50.6 33.6 24.3
f. GROVE - FT (Ours) 21.0 77.7 15.8 18.1
g. GROVE - PT+FT (Ours) 21.4 83.5 40.0 28.7
@@
GROVE PT+FT (Ours) 63.7
Table 3: State-of-the-art comparison of spatial grounding on the VidSTG [48] test set (declarative sentences). All models use ground truth temporal localization. Large-scale pretraining (PT+FT) results in an improvement over fine-tuning only (FT) for our model GROVE.
Method FT $\text{m}_{\text{sIoU}}$
PG-V-L (13B) [25] 35.1
GLaMM [31]
+ SAM2 [32] 38.6
GROVE 43.0
VideoGLaMM[26] 39.7
GROVE 55.5
Table 4: State-of-the-art comparison of spatial grounding on the VidSTG [48] test set (interrogative sentences). All models use ground truth temporal localization. GROVE outperforms all competitors both in a pre-training only setting (✗) and when fine-tuned on VidSTG (✓).
@@
Figure 4: Qualitative examples showing predictions of our GROVE model. Please note that GROVE is able to: (i) produce video-level natural language captions describing the main action in the video; (ii) ground multiple objects; and (iii) produce spatio-temporally consistent bounding box predictions. Please note that the second row shows an example of model prediction that is partly incorrect as the blue box, while temporally consistent, does not depict a “yarn”. More results including failure modes are discussed in appendix A.